# sample.bib

### v2

@@ 3.3 Inter-slice Fusion Transformer
In this section, inter-slice fusion transformer (IF-Trans) is proposed to capture inter-slice cues, as shown in Fig. 3. We utilize IF-Trans at multi-scales and discuss the benefit of multi-scale learning in Section 5.3. Inputs of the $k$-th IF-Trans are feature maps $\left\{\bm{f}_{i-s}^{k},\cdots,\bm{f}_{i}^{k},\cdots,\bm{f}_{i+s}^{k}\right\}$ extracted by the encoder and the momentum encoder, where $k$ represents the $k$-st scale of two encoders ($k=1,2,3,4$). The neighborhood slice number $s$ is set to $1$ in our default configuration. We give a detailed analysis of $s$ in Section 5.3. Therefore, the model uses adjacent ($1$-th neighborhood) slices of the target slice $\bm{x}_{i}$ as additional inputs.
Figure 3: Schematic of inter-slice fusion transformer (IF-Trans) module. The neighborhood slice number is set to $1$ in this figure, consistent with our default model configuration. It has two successive IF-Trans with different window partitioning configurations. The window-based self-attention is expanded to the inter-slice dimension, promoting target slice pixels to learn intra- and inter-slice contexts.
Different from standard self-attention [Vaswani2017attn] with quadratic complexity, the proposed IF-Trans only calculates self-attention within the local window. As shown in the left part of Fig. 3, feature maps are portioned to several non-overlapping windows [^1] For intuitive explanation, feature maps are replaced by input images, and the number of pixels is simplified to 16.. Compared with Swin transformer [liu2021swin], we compute self-attention within inter- and intra-slice local windows (i.e., CSW-MSA, cross-slice window-based multi-head self-attention) instead of intra-slice local windows. Therefore, the orange pixel in the target slice can capture not only intra-slice information (yellow pixels) but also perceives inter-slice information (green pixels).
However, the local CSW-MSA lacks connections across windows, degrading its feature formulation power. Similar to [liu2021swin], a shifted window partitioning strategy is introduced, allowing each pixel to receive broader views from intra- and inter-slices. In Fig. 3, the first transformer module adopts a regular window partition approach, and the feature map is evenly divided into $2\times 2$ windows of size $2\times 2$ ($M=2$) [^2] To correspond with Fig. 3, $M$ is set to 2 here.. The second transformer module uses a different partitioning configuration. Windows of the preceding layer are displaced by $\left(\lfloor\frac{M}{2}\rfloor,\lfloor\frac{M}{2}\rfloor\right)$ pixels to generate new windows. By doing so, the orange pixel can conduct self-attention (i.e., SCSW-MSA, a shifted version of CSW-MSA) with more pixels, boosting representation abilities. In practice, these two configurations are served as two consecutive layers to get an IF-Trans module. Outputs of IF-Trans can be formulated as:
@@
where $\hat{\bm{f}}^{k,l}_{i}$ and $\bm{f}^{k,l}_{i}$ represents output feature maps of the (S)CSW-MSA module $\mathcal{T}^{\rm(S)}$ and the multilayer perceptron (MLP) module $\mathcal{M}$ in the $l$-th layer, respectively. LN indicates layer normalization. The query-key-value (QKV) self-attrntion [Vaswani2017attn] in (S)CSW-MSA is computed as follows:
@@ 3.4 Loss Function
Following previous methods [xie2021cotr][hatamizadeh2022unetr][zhou2023nnformer], our model is trained end-to-end using the deep supervision strategy [lee2015deeply]. As illustrated in Fig. 2, final segmentation results are generated by the segmentation head ($1\times 1$ convolutional layer). Additionally, two smaller resolutions of decoder outputs are selected as auxiliary supervision signals. The deep supervision path in Fig. 2 consists of an upsample layer and a $1\times 1$ convolutional layer. Therefore, the loss function can be formulated as follows:
@@ 4.1 Datasets
To thoroughly compare MOSformer to previous methods, we conduct experiments on three challenging benchmarks: the Synapse multi-organ segmentation dataset [landman2015miccai], the automated cardiac diagnosis challenge (ACDC) dataset [bernard2018deep], and the abdominal organ segmentation (AMOS) dataset [ji2022amos].
Synapse for Multi-organ Segmentation. This dataset consists of 30 abdominal CT scans with 8 organs (aorta, gallbladder, left kidney, right kidney, liver, pancreas, spleen, and stomach). Each volume has $85\sim 198$ slices of $512\times 512$ pixels. Following the splits adopted in TransUNet [chen2024transunet], the dataset is divided into 18 training cases and 12 testing cases.
ACDC for Automated Cardiac Diagnosis Challenge. The ACDC dataset includes cardiac MRI images of 100 patients from real clinical exams with manual annotations of left ventricle (LV), right ventricle (RV), and myocardium (Myo). Consistent with TransUNet [chen2024transunet], the dataset is split into 70 training cases, 10 validation cases, and 20 testing cases.
@@ 4.2 Implementation Details
All experiments are implemented based on PyTorch 1.12.0, Python 3.8, and Ubuntu 18.04. Our model is trained on a single Nvidia A6000 GPU with 48GB of memory. The same model configurations are utilized on three datasets. Input medical images are resized to $224\times 224$ for fair comparison. SGD optimizer with momentum of $0.9$ and weight decay of $1e^{-4}$ is adopted to train our model for 300 epochs. The batch size is set to 24. A cosine learning rate scheduler with five epochs of linear warm-up is used during training, and the maximum and minimum learning rates are $3e^{-2}$ and $5e^{-3}$, respectively.
@@ 5.1 Comparisons with SOTAs
Table 1: Comparison with state-of-the-art models on the multi-organ segmentation (Synapse) dataset. The best results are highlighted in blue and the second-best results are highlighted in red. The evaluation metrics are DSC and HD95, consisting with TransUNet [chen2024transunet]. Moreover, DSC of each organ is reported in this table. means the results are borrowed from [zhou2023nnformer].
Dimension Method Average Aotra Gallbladder Kidney (L) Kidney (R) Liver Pancreas Spleen Stomach
@@
2D UNet [ronneberger2015u] [MICCAI’15] 76.85 39.70 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
AttnUNet [schlemper2019attention] [MedIA’19] 77.77 36.02 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75
TransUNet [chen2024transunet] [MedIA’24] 77.48 31.69 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62
MISSFormer [huang2023missformer] [TMI’23] 81.96 18.20 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
SwinUNet [cao2022swin] [ECCVW’22] 79.12 21.55 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
MT-UNet [wang2022mixed] [ICASSP’22] 78.59 26.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
UCTransNet [wang2022uctransnet] [AAAI’22] 78.23 26.75 88.86 66.97 80.19 73.18 93.17 56.22 87.84 79.43
CASTformer [you2022class] [NeurIPS’22] 82.55 22.73 89.05 67.48 86.05 82.17 95.61 67.49 91.00 81.55
HiFormer [heidari2023hiformer] [WACV’23] 80.39 14.70 86.21 65.69 85.23 79.77 94.61 59.52 90.99 81.08
3D V-Net [milletari2016v] [3DV’16] 68.81 - 75.34 51.87 77.10 80.75 87.84 40.05 80.56 56.98
CoTr [xie2021cotr] [MICCAI’21] 80.78 19.15 85.42 68.93 85.45 83.62 93.89 63.77 88.58 76.23
UNETR [hatamizadeh2022unetr] [WACV’22] 79.56 22.97 89.99 60.56 85.66 84.80 94.46 59.25 87.81 73.99
SwinUNETR [hatamizadeh2021swin] [MICCAIW’22] 83.51 14.78 90.75 66.72 86.51 85.88 95.33 70.07 94.59 78.20
nnFormer [zhou2023nnformer] [TIP’23] 86.57 10.63 92.04 70.17 86.57 86.25 96.84 83.35 90.51 86.83
2.5D AFTer-UNet [yan2022after] [WACV’22] 81.02 - 90.91 64.81 87.90 85.30 92.20 63.54 90.99 72.48
TransUNet-2.5D [zhang2023multi] [TIM’23] 84.24 19.24 89.97 73.28 83.99 81.33 95.61 70.39 94.59 84.78
@@
Figure 4: Visual comparisons with current state-of-the-art methods on the multi-organ segmentation (Synapse) dataset.
Multi-organ Segmentation (Synapse). Quantitative results of state-of-the-art models and our MOSformer are presented in Table 1. MOSformer achieves $85.63\%$ DSC and $13.40$ ${\rm mm}$ HD95 on this dataset. Compared with the best 2D-based methods, i.e., CASTformer [you2022class], MOSformer is able to surpass it by a large margin ($+3.08\%$ DSC and $-9.33$ ${\rm mm}$ HD95). For 2.5D-based baselines, MOSformer demonstrates notable performance enhances, offers at least $+4.61\%$ and $+1.39\%$ DSC gains over AFTer-UNet [yan2022after] and TransUNet-2.5D [zhang2023multi]. These results indicate i) the necessity of inter-slice information in 3D medical image segmentation; and ii) the effectiveness of distinguishable and consistent slice features produces by dual encoders with a momentum update.
Table 2: Comparison with the state-of-the-art models on the automated cardiac diagnosis challenge (ACDC) dataset. The best results are highlighted in blue and the second-best results are highlighted in red. We only report DSC in this table, following the evaluation setting of TransUNet [chen2024transunet]. Moreover, DSC of each anatomical structure is reported in this table. means the results are borrowed from [zhou2023nnformer]. means the baselines are implemented by ourselves.
Dimension Method Average DSC (%) $\uparrow$ RV Myo LV
2D UNet [ronneberger2015u] [MICCAI’15] 87.60 84.62 84.52 93.68
AttnUNet [schlemper2019attention] [MedIA’19] 86.90 83.27 84.33 93.53
TransUNet [chen2024transunet] [MedIA’24] 89.71 86.67 87.27 95.18
MISSFormer [huang2023missformer] [TMI’23] 91.19 89.85 88.38 95.34
SwinUNet [cao2022swin] [ECCVW’22] 88.07 85.77 84.42 94.03
MT-UNet [wang2022mixed] [ICASSP’22] 90.43 86.64 89.04 95.62
UCTransNet [wang2022uctransnet] [AAAI’22] 91.98 90.06 89.87 96.02
HiFormer [heidari2023hiformer] [WACV’23] 90.40 88.24 87.63 95.30
3D UNETR [hatamizadeh2022unetr] [WACV’22] 88.61 85.29 86.52 94.02
nnFormer [zhou2023nnformer] [TIP’23] 92.06 90.94 89.58 95.65
2.5D CAT-Net [hung2022cat] [TMI’22] 90.02 86.05 88.75 95.27
@@
Figure 5: Visual comparisons with current state-of-the-art methods on the automatic cardiac diagnosis challenge (ACDC) dataset.
We also compare MOSformer with 3D-based segmentation methods. MOSformer still has competitive performance, surpassing four of the most widely recognized models and achieving comparable performance to nnFormer [zhou2023nnformer]. It should be noted that MOSformer obtains better DSC than nnFormer in four organs (half of the categories), including gallbladder ($+1.73\%$), left kidney ($+3.75\%$), spleen ($+1.78\%$), and stomach ($+1.04\%$). Among these organs, gallbladder and stomach are two of the most difficult organs to segment since gallbladder is very small and boundaries between gallbladder and liver are blurred while stomach has a significant intra-class variance. This reveals that our MOSformer can learn more discriminative features and have comprehensive understandings of organ structures.
Table 3: Comparison with the state-of-the-art models on the abdominal organ segmentation (AMOS) dataset. The best results are highlighted in blue and the second-best results are highlighted in red. DSC is utilized as evaluation metric. Moreover, DSC of each organ is reported in this table. means the baselines are implemented by ourselves.
Dimension Method Average DSC (%) $\uparrow$ Spleen Kid. (R) Kid. (L) Gall. Eso. Liver Stom. Aotra IVC Panc. Adr. (R) Adr. (L) Duo. Blad. Pros.
2D UNet [ronneberger2015u] [MICCAI’15] 82.53 92.25 92.45 92.50 81.85 79.98 94.73 84.80 92.20 82.94 77.35 67.13 69.34 72.77 82.40 75.31
TransUNet [chen2024transunet] [MedIA’24] 80.10 91.26 92.47 91.90 78.01 77.00 94.93 80.04 91.98 82.99 74.30 63.66 53.84 71.65 81.37 76.03
MISSFormer [huang2023missformer] [TMI’23] 78.16 93.13 91.98 91.88 75.89 71.87 94.27 80.14 88.74 77.53 71.39 60.65 59.32 64.43 77.97 73.16
UCTransNet [wang2022uctransnet] [AAAI’22] 82.34 93.37 92.32 91.90 77.09 79.77 94.78 85.95 91.77 82.84 77.44 65.88 68.98 71.36 83.93 77.71
HiFormer [heidari2023hiformer] [WACV’23] 80.03 92.73 92.79 92.01 79.44 76.42 94.55 82.65 90.56 80.16 73.59 61.14 58.73 68.12 82.01 75.64
3D UNETR [hatamizadeh2022unetr] [WACV’22] 78.07 93.38 93.00 92.28 73.17 69.72 94.86 73.25 90.82 80.20 73.44 65.19 60.69 65.46 74.10 71.49
nnFormer [zhou2023nnformer] [TIP’23] 78.66 91.43 92.39 92.08 76.74 69.16 94.95 84.84 89.53 82.06 75.91 62.56 60.36 68.50 74.74 64.61
2.5D MOSformer [Ours] 85.43 95.26 94.68 94.54 81.53 82.05 96.55 89.07 92.81 86.16 80.28 73.28 73.19 75.05 86.92 80.05
Figure 6: Visual comparisons with current state-of-the-art methods on the abdominal organ segmentation (AMOS) dataset.
@@
Automated Cardiac Diagnosis Challenge (ACDC). To further prove model generalizing performance, MOSformer is evaluated on the automated cardiac diagnosis challenge (ACDC) dataset. It should be noted that MRI images in this dataset can be considered anisotropic since they have high in-plane image resolution (e.g., $1.37\sim 1.68$ ${\rm mm}$) and low throughplane resolution (e.g., $5$ ${\rm mm}$[bernard2018deep]. Quantitative results are summrized in Table 2. Compared with state-of-the-art methods (2D, 2.5D, and 3D-based), MOSformer achieves the best performance with $92.19\%$ DSC. Thus, the above results indicate that our 2.5D-based MOSformer is more advantageous to process anisotropic data compared with 3D-based models. Fig. 5 presents qualitative comparisons for different methods on this dataset. As seen, MOSformer can locate anatomical structures more accurately. Specifically, in case 3, many models mistakenly classify regions outside the myocardium into the right ventricle while MOSformer donot produce any false positive predictions.
@@
Abdominal Organ Segmentation (AMOS). Additionally, a large dataset with 200 training cases and 100 testing cases is also adopted in our experiments. Overall results and individual DSC on 15 organs are reported, as shown in Table 3. Our MOSformer achieves the best DSC in 14 organs and the second-best DSC in one organ. In a surprise, MOSformer offers $6.77\%$ DSC improvement over 3D-based nnFormer while they have close performance on the multi-organ segmentation (Synapse) dataset. Based on the above observation, it can be concluded that the performance of MOSformer is more stable across different datasets compared with nnFormer. Visualization results are shown in Fig. 6. Compared with baselines, our MOSformer is able to accurately segment organs of diverse shapes and sizes, thus providing more consistent results with ground truth.
@@
Table 4: Ablation studies of each component on the multi-organ segmentation (Synapse) and the automated cardiac diagnosis challenge (ACDC) datasets. Enc-S: Single encoder; Enc-D: Dual encoders; Enc-DM: Dual encoders with a momentum update. The best results are highlighted in bold. means the model is 2D-based.
Model Module Synapse ACDC
Enc-S Enc-D Enc-DM IF-Trans DSC (%) $\uparrow$ DSC (%) $\uparrow$
Model-1 82.42 (-3.21) 91.61 (-0.58)
Model-2 84.23 (-1.40) 92.04 (-0.15)
Model-3 84.93 (-0.70) 92.10 (-0.09)
MOSformer 85.63 92.19

### v3

@@ 3.3 Inter-slice Fusion Transformer
In this section, inter-slice fusion transformer (IF-Trans) is proposed to capture inter-slice cues, as shown in Fig. 3. We utilize IF-Trans at multiple scales and discuss the benefit of multi-scale learning in Section 5.3. Inputs of the $k$-th IF-Trans are feature maps $\left\{\bm{f}_{i-s}^{k},\cdots,\bm{f}_{i}^{k},\cdots,\bm{f}_{i+s}^{k}\right\}$ extracted by the encoder and the momentum encoder, where $k$ represents the $k$-th scale of two encoders ($k=1,2,3,4$). The neighborhood slice number $s$ is set to $1$ in our default configuration. We give a detailed analysis of $s$ in Section 5.3. Therefore, the model uses adjacent ($1$-st neighborhood) slices of the target slice $\bm{x}_{i}$ as additional inputs.
Different from standard self-attention [49] with quadratic complexity, the proposed IF-Trans only calculates self-attention within the local window. As shown in the left part of Fig. 3, feature maps are partitioned into several non-overlapping windows [^1] For intuitive explanation, feature maps are replaced by input images, and the number of feature pixels is simplified to 16.. Compared with Swin transformer [30], we compute self-attention within inter- and intra-slice local windows (i.e., CSW-MSA, cross-slice window-based multi-head self-attention) instead of intra-slice local windows. Consequently, the purple feature pixel in the target slice attends not only to the yellow pixels within its own slice but also to the green pixels from neighborhood slices, thereby capturing both intra-slice and inter-slice context, as illustrated by the red arrows in Fig. 3.
However, the local CSW-MSA lacks connections across windows, reducing its representational power. Similar to [30], a shifted window partitioning strategy is introduced, allowing each pixel to receive broader views from intra- and inter-slices. In Fig. 3, the first transformer module adopts a regular window partition approach, and the feature map is evenly divided into $2\times 2$ windows of size $2\times 2$ (M=2M=2) [^2] To correspond with Fig. 3, $M$ is set to 2 here.. The second transformer module uses a different partitioning configuration. Windows of the preceding layer are displaced by $\left(\lfloor\frac{M}{2}\rfloor,\lfloor\frac{M}{2}\rfloor\right)$ pixels to generate new windows. By doing so, the orange pixel can conduct self-attention (i.e., SCSW-MSA, a shifted version of CSW-MSA) with more pixels, thereby boosting its representational capacity. In practice, these two configurations are served as two consecutive layers to get an IF-Trans module. Outputs of IF-Trans can be formulated as:
@@
where $\hat{\bm{f}}^{k,l}_{i}$ and $\bm{f}^{k,l}_{i}$ represents output feature maps of the (S)CSW-MSA module $\mathcal{T}^{\rm(S)}$ and the multilayer perceptron (MLP) module $\mathcal{M}$ in the $l$-th layer, respectively. LN indicates layer normalization. The query-key-value (QKV) self-attention [49] in (S)CSW-MSA is computed as follows:
@@ 3.4 Loss Function
Following previous methods [45][44][46], our model is trained end-to-end using the deep supervision strategy [50]. As illustrated in Fig. 2, final segmentation results are generated by the segmentation head ($1\times 1$ convolutional layer). Additionally, two smaller resolutions of decoder outputs are selected as auxiliary supervision signals. The deep supervision path in Fig. 2 consists of an upsample layer and a $1\times 1$ convolutional layer. Therefore, the loss function can be formulated as follows:
@@ 4.1 Datasets
To thoroughly compare MOSformer to previous methods, we conduct experiments on three challenging benchmarks: the Synapse multi-organ segmentation dataset [51], the automated cardiac diagnosis challenge (ACDC) dataset [52], and the abdominal organ segmentation (AMOS) dataset [53].
Synapse for Multi-organ Segmentation. This dataset consists of 30 abdominal CT scans with 8 organs (aorta, gallbladder, left kidney, right kidney, liver, pancreas, spleen, and stomach). Each volume has $85\sim 198$ slices of $512\times 512$ pixels. Following the splits adopted in TransUNet [23], the dataset is divided into 18 training cases and 12 testing cases.
ACDC for Automated Cardiac Diagnosis Challenge. The ACDC dataset includes cardiac MRI images of 100 patients from real clinical exams with manual annotations of left ventricle (LV), right ventricle (RV), and myocardium (Myo). Consistent with TransUNet [23], the dataset is split into 70 training cases, 10 validation cases, and 20 testing cases.
@@ 4.2 Implementation Details
All experiments are implemented based on PyTorch 1.12.0, Python 3.8, and Ubuntu 18.04. Our model is trained on a single Nvidia A6000 GPU with 48GB of memory. The same model configurations are utilized on three datasets. Input medical images are resized to $224\times 224$ for a fair comparison. SGD optimizer with momentum of $0.9$ and weight decay of $1e^{-4}$ is adopted to train our model for 300 epochs. The batch size is set to 24. A cosine learning rate scheduler with five epochs of linear warm-up is used during training, and the maximum and minimum learning rates are $3e^{-2}$ and $5e^{-3}$, respectively.
@@ 5.1 Comparisons with SOTAs
Table 1: Comparison with state-of-the-art models on the multi-organ segmentation (Synapse) dataset. The best results are highlighted in bold and the second-best results are underlined. The evaluation metrics are DSC and HD95, consistent with TransUNet [23]. Moreover, DSC of each organ is reported in this table. and indicate the results are borrowed from [46] and [22], respectively. means the baselines are implemented by ourselves. Baselines without any symbol represent the results are from the original papers.
Dimension Method Average Aorta Gallbladder Kidney (L) Kidney (R) Liver Pancreas Spleen Stomach
@@
2D UNet [12] [MICCAI’15] 76.85 39.70 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
AttnUNet [54] [MedIA’19] 77.77 36.02 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75
TransUNet [23] [MedIA’24] 77.48 31.69 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62
MISSFormer [43] [TMI’23] 81.96 18.20 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
SwinUNet [22] [ECCVW’22] 79.12 21.55 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
MT-UNet [55] [ICASSP’22] 78.59 26.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
UCTransNet [56] [AAAI’22] 78.23 26.75 88.86 66.97 80.19 73.18 93.17 56.22 87.84 79.43
CASTformer [42] [NeurIPS’22] 82.55 22.73 89.05 67.48 86.05 82.17 95.61 67.49 91.00 81.55
HiFormer [57] [WACV’23] 80.39 14.70 86.21 65.69 85.23 79.77 94.61 59.52 90.99 81.08
SAMed [58] [arXiv’23] 81.88 20.64 87.77 69.11 80.45 79.95 94.80 72.17 88.72 82.06
3D V-Net [59] [3DV’16] 68.81 - 75.34 51.87 77.10 80.75 87.84 40.05 80.56 56.98
CoTr [45] [MICCAI’21] 80.78 19.15 85.42 68.93 85.45 83.62 93.89 63.77 88.58 76.23
UNETR [44] [WACV’22] 79.56 22.97 89.99 60.56 85.66 84.80 94.46 59.25 87.81 73.99
SwinUNETR [60] [MICCAIW’22] 83.51 14.78 90.75 66.72 86.51 85.88 95.33 70.07 94.59 78.20
nnFormer [46] [TIP’23] 86.57 10.63 92.04 70.17 86.57 86.25 96.84 83.35 90.51 86.83
SAM3D [61] [ISBI’24] 79.56 17.87 89.57 49.81 86.31 85.64 95.42 69.32 84.29 76.11
2.5D AFTer-UNet [27] [WACV’22] 81.02 - 90.91 64.81 87.90 85.30 92.20 63.54 90.99 72.48
TransUNet-2.5D [62] [TIM’23] 84.24 19.24 89.97 73.28 83.99 81.33 95.61 70.39 94.59 84.78
CSA-Net [38] [CIBM’24] 79.96 32.11 83.91 64.99 83.56 79.93 94.43 62.65 91.12 79.12
@@
We select several state-of-the-art 2D, 3D, and 2.5D medical image segmentation models as our baselines. To ensure a fair comparison, all models are trained and evaluated using identical preprocessing pipelines. Specifically, we apply the TransUNet preprocessing protocol [23] to both the multi-organ segmentation (Synapse) and the automated cardiac diagnosis challenge (ACDC) datasets, and follow the preprocessing procedures of [53] for the abdominal organ segmentation (AMOS) dataset. It should be noted that we only visualize selected qualitative results from some representative models for clarity and visual impact.
Multi-organ Segmentation (Synapse). Quantitative results of state-of-the-art models and our MOSformer are presented in Table 1. MOSformer achieves $85.63\%$ DSC and $13.40$ $\rm mm$} HD95 on this dataset. Compared with the best 2D-based method, i.e., CASTformer [42], MOSformer is able to surpass it by a large margin ($3.08\%$ DSC and 9.33-9.33 $\rm mm$} HD95). For 2.5D-based baselines, MOSformer demonstrates notable performance enhancements, offering at least $4.61\%$, $1.39\%$, and $5.67\%$ DSC gains over AFTer-UNet [27], TransUNet-2.5D [62], and CSA-Net [38], respectively. These results indicate i) the necessity of inter-slice information in 3D medical image segmentation; and ii) the effectiveness of distinguishable and consistent slice features produced by dual encoders with a momentum update.
We also compare our MOSformer with 3D-based segmentation methods. It still has competitive performance, surpassing five of the most widely recognized models and achieving comparable performance to nnFormer [46]. It should be noted that MOSformer obtains better DSC than nnFormer in four organs (half of the categories), including gallbladder ($1.73\%$), left kidney ($3.75\%$), spleen ($1.78\%$), and stomach ($1.04\%$). Among these organs, gallbladder and stomach are two of the most difficult organs to segment since the gallbladder is very small and the boundaries between the gallbladder and the liver are blurred while the stomach has a significant intra-class variance. This reveals that our MOSformer can learn more discriminative features and has a comprehensive understanding of organ structures.
Figure 4: Visual comparisons with some representative methods on the multi-organ segmentation (Synapse) dataset.
Table 2: Comparison with the state-of-the-art models on the automated cardiac diagnosis challenge (ACDC) dataset. The best results are highlighted in bold and the second-best results are underlined. We only report DSC in this table, following the evaluation setting of TransUNet [23]. Moreover, DSC of each anatomical structure is reported in this table. and indicate the results are borrowed from [46] and [55], respectively. means the baselines are implemented by ourselves. Baselines without any symbol represent the results are from the original papers.
Dimension Method DSC (%) $\uparrow$ RV Myo LV
2D UNet [12] [MICCAI’15] 87.60 84.62 84.52 93.68
AttnUNet [54] [MedIA’19] 86.90 83.27 84.33 93.53
TransUNet [23] [MedIA’24] 89.71 86.67 87.27 95.18
MISSFormer [43] [TMI’23] 91.19 89.85 88.38 95.34
SwinUNet [22] [ECCVW’22] 88.07 85.77 84.42 94.03
MT-UNet [55] [ICASSP’22] 90.43 86.64 89.04 95.62
UCTransNet [56] [AAAI’22] 91.98 90.06 89.87 96.02
HiFormer [57] [WACV’23] 90.40 88.24 87.63 95.30
3D UNETR [44] [WACV’22] 88.61 85.29 86.52 94.02
nnFormer [46] [TIP’23] 92.06 90.94 89.58 95.65
SAM3D [61] [ISBI’24] 90.41 89.44 87.12 94.67
2.5D CAT-Net [28] [TMI’22] 90.02 86.05 88.75 95.27
CSA-Net [38] [CIBM’24] 89.58 86.56 86.91 95.26
@@
Figure 5: Visual comparisons with some representative methods on the automatic cardiac diagnosis challenge (ACDC) dataset.
@@
Automated Cardiac Diagnosis Challenge (ACDC). To further prove the model’s generalization performance, MOSformer is evaluated on the automated cardiac diagnosis challenge (ACDC) dataset. It should be noted that MRI images in this dataset can be considered anisotropic since they have high in-plane image resolution (e.g., $1.37\sim 1.68$ $\rm mm$}) and low through-plane resolution (e.g., 55 $\rm mm$}[52]. Quantitative results are summarized in Table 2. Compared with state-of-the-art methods (2D, 2.5D, and 3D-based), MOSformer achieves the best performance with $92.19\%$ DSC. Thus, the above results indicate that our 2.5D-based MOSformer is more effective at processing anisotropic data compared with 3D-based models. Fig. 5 presents qualitative comparisons for different methods on this dataset. As seen, MOSformer can locate anatomical structures more accurately. Specifically, in case 3, many models mistakenly classify regions outside the myocardium into the right ventricle while MOSformer does not produce any false positive predictions.
Table 3: Comparison with the state-of-the-art models on the abdominal organ segmentation (AMOS) dataset. The best results are highlighted in bold and the second-best results are underlined. DSC is utilized as evaluation metric. Moreover, DSC of each organ is reported in this table. means the baselines are implemented by ourselves.
Dimension Method DSC (%) $\uparrow$ Spleen Kid. (R) Kid. (L) Gall. Eso. Liver Stom. Aorta IVC Panc. Adr. (R) Adr. (L) Duo. Blad. Pros.
2D UNet [12] [MICCAI’15] 82.53 92.25 92.45 92.50 81.85 79.98 94.73 84.80 92.20 82.94 77.35 67.13 69.34 72.77 82.40 75.31
TransUNet [23] [MedIA’24] 80.10 91.26 92.47 91.90 78.01 77.00 94.93 80.04 91.98 82.99 74.30 63.66 53.84 71.65 81.37 76.03
MISSFormer [43] [TMI’23] 78.16 93.13 91.98 91.88 75.89 71.87 94.27 80.14 88.74 77.53 71.39 60.65 59.32 64.43 77.97 73.16
UCTransNet [56] [AAAI’22] 82.34 93.37 92.32 91.90 77.09 79.77 94.78 85.95 91.77 82.84 77.44 65.88 68.98 71.36 83.93 77.71
HiFormer [57] [WACV’23] 80.03 92.73 92.79 92.01 79.44 76.42 94.55 82.65 90.56 80.16 73.59 61.14 58.73 68.12 82.01 75.64
3D UNETR [44] [WACV’22] 78.07 93.38 93.00 92.28 73.17 69.72 94.86 73.25 90.82 80.20 73.44 65.19 60.69 65.46 74.10 71.49
nnFormer [46] [TIP’23] 78.66 91.43 92.39 92.08 76.74 69.16 94.95 84.84 89.53 82.06 75.91 62.56 60.36 68.50 74.74 64.61
2.5D CSA-Net [38] [CIBM’24] 82.12 91.25 93.51 93.68 79.01 78.80 95.32 82.14 91.64 83.94 75.18 68.27 69.37 71.36 83.00 75.33
MOSformer [Ours] 85.43 95.26 94.68 94.54 81.53 82.05 96.55 89.07 92.81 86.16 80.28 73.28 73.19 75.05 86.92 80.05
Figure 6: Visual comparisons with some representative methods on the abdominal organ segmentation (AMOS) dataset.
@@
Abdominal Organ Segmentation (AMOS). Additionally, a large dataset with 200 training cases and 100 testing cases is also adopted in our experiments. Overall results and individual DSC on 15 organs are reported, as shown in Table 3. Our MOSformer achieves the best DSC in 14 organs and the second-best DSC in one organ. Surprisingly, MOSformer offers $6.77\%$ DSC improvement over 3D-based nnFormer while they have similar performance on the multi-organ segmentation (Synapse) dataset. Based on the above observation, it can be concluded that the performance of MOSformer is more stable across different datasets compared with nnFormer. Visualization results are shown in Fig. 6. Compared with baselines, our MOSformer is able to accurately segment organs of diverse shapes and sizes, thus providing more consistent results with ground truths.
Statistical Analysis. In Table 4, we report Wilcoxon-test $p$-values on case-level DSC for paired comparisons between MOSformer and the strongest publicly available 2D, 3D, and 2.5D baselines. On the Synapse dataset, although the 3D nnFormer attains a slightly higher mean DSC than MOSformer, the difference is not statistically significant ($p$ = 0.339), indicating comparable performance. In contrast, MOSformer significantly outperforms the 2D MISSFormer ($p$ < 0.001) and the 2.5D CSA-Net ($p$ < 0.001). On the ACDC dataset, MOSformer achieves the highest mean DSC, but the margins over UCTransNet ($p$ = 0.745) and nnFormer ($p$ = 0.826) are not significant, while the improvement over CAT-Net is significant ($p$ < 0.001). On the AMOS dataset, MOSformer shows significant gains over all baselines (all $p$ < 0.001), demonstrating consistent advantages. Notably, the AMOS dataset has a much larger test set ($N$ = 100) than the Synapse ($N$ = 12) and the ACDC ($N$ = 40) datasets, providing greater statistical power. Accordingly, $p$-values on the AMOS dataset are more sensitive to performance differences, whereas non-significant results on the Synapse or the ACDC datasets may reflect limited sample sizes. Overall, this analysis strengthens the empirical evidence for MOSformer’s effectiveness across diverse datasets.
Table 4: Statistical analysis on the multi-organ segmentation (Synapse), automated cardiac diagnosis challenge (ACDC), and abdominal organ segmentation (AMOS) datasets. NN is the number of testing cases.
Synapse ($N$ = 12) ACDC ($N$ = 40) AMOS ($N$ = 100)
Method DSC $p$ Method DSC $p$ Method DSC $p$
MISSFormer [43] [TMI’23] 81.96 < 0.001 UCTransNet [56] [AAAI’22] 91.89 0.745 UNet [12] [MICCAI’15] 82.53 < 0.001
nnFormer [46] [TIP’23] 86.57 0.339 nnFormer [46] [TIP’23] 92.06 0.826 nnFormer [46] [TIP’23] 78.66 < 0.001
CSA-Net [38] [CIBM’24] 79.96 < 0.001 CAT-Net [28] [TMI’22] 90.02 < 0.001 CSA-Net [38] [CIBM’24] 82.12 < 0.001
MOSformer [Ours] 85.63 - MOSformer [Ours] 92.19 - MOSformer [Ours] 85.43 -
@@
Table 5: Ablation study of each component on the multi-organ segmentation (Synapse), automated cardiac diagnosis challenge (ACDC), and abdominal organ segmentation (AMOS) datasets. Enc-S: Single encoder; Enc-D: Dual encoders; Enc-DM: Dual encoders with a momentum update. The best results are highlighted in bold. means the model is 2D-based.
Model Module Synapse ACDC AMOS
Enc-S Enc-D Enc-DM IF-Trans DSC (%) $\uparrow$ DSC (%) $\uparrow$ DSC (%) $\uparrow$
Model-1 82.42 (-3.21) 91.61 (-0.58) 81.28 (-4.15)
Model-2 84.23 (-1.40) 92.04 (-0.15) 82.63 (-2.80)
Model-3 84.93 (-0.70) 92.10 (-0.09) 83.88 (-1.55)
MOSformer 85.63 92.19 85.43